# Language Bottleneck Models: A Framework for Interpretable Knowledge Tracing and Beyond

### v1

@@ 3.3 Training Objective and Optimization
To optimize $f_{\theta}$, we employ Group Relative Policy Optimization (GRPO) shao2024deepseekmath . For each input $X$, the encoder generates $G$ candidate summaries $\{S_{1},\dots,S_{G}\}$, each evaluated using $\mathcal{R}(S_{i};g)$. We compute group-relative advantage and update parameters:
$A(S_{i})=\frac{\mathcal{R}(S_{i};g)-\mu}{\sigma},\quad\nabla_{\theta}\mathbb{E}_{S\sim f_{\theta}(X)}[\mathcal{R}(S;g)]\approx\frac{1}{G}\sum_{i=1}^{G}A(S_{i})\cdot\nabla_{\theta}\log p_{\theta}(S_{i}|X),$ (2)
where $\mu$ and $\sigma$ are the mean and standard deviation of rewards within the group. This approach encourages the encoder to produce summaries that maximize the reward function defined in Eq. 1.
3.4 Steerability of the Estimated Knowledge States
A key advantage of LBMs is the ability for humans to steer the model’s behavior. We identify three mechanisms for human-model interaction (detailed in Appendix):
  • Prompt engineering the encoder. Directly shaping summary generation through input instructions.
  • Steering via reward signals. Incorporating human preferences during encoder training through the $\Omega(S)$ term in Eq. 1.
@@ 3.4 Steerability of the Estimated Knowledge States
  • Augmenting with student-specific information. Educators can supplement the model with additional knowledge not present in observed data—either by augmenting encoder inputs or by editing the generated summary before decoding. This enables integration of recent classroom observations or specific misconceptions identified through in-person interactions.
@@ 4 Related Work
In this section, we review the literature related to our proposed approach, spanning traditional knowledge tracing, recent advances in LLM-based student modeling, and concept bottleneck models (CBMs), and text summarization models.
Traditional Knowledge Tracing
Traditional knowledge tracing (KT) methods have long been used to model student learning and predict future performance shen2024survey . Early works such as Bayesian Knowledge Tracing (BKT) corbett_knowledge_1994 model student mastery as a probabilistic process, while more recent methods like Deep Knowledge Tracing (DKT) piech_deep_2015 and Attentive Knowledge Tracing (AKT) ghosh2020context uses an attention-based architecture to learn contextual representation of questions and student knowledge states. Despite their strong predictive performance, these models represent a student’s knowledge state as an abstract, high-dimensional latent vector, which poses significant challenges in interpretability and actionable feedback for educators.
LLM-Based Knowledge Tracing
Recent studies have begun integrating Large Language Models (LLMs) into the KT framework. For example, li2024explainable demonstrated that LLMs are able to make sensible predictions about student responses when prompted with adequate information. Other works lee2024language ; kim2024beyond have studied how LLMs can help mitigate the cold-start problem compared to traditional KT approaches, while wang2025llm demonstrated state-of-the-art performances in KT by combining LLMs with sequence interaction models. However, these methods generally remain opaque: they either treat the LLM as a black-box, or rely on model-generated explanations that are susceptible to hallucination (bender2021dangers, ).
Concept Bottleneck Models
@@
Concept Bottleneck Models (CBMs) koh2020concept improve interpretability through human-understandable concept activations as intermediates, with extensions exploring unsupervised concept learning oikarinen2023label , test-time interventions shin2023closer , and theoretical analyses of concept set design luyten2024theoretical . However, CBMs typically rely on finite predefined concept sets, limiting their applicability to complex tasks like knowledge tracing. Recently, yamaguchitoward introduced Explanation Bottleneck Models (XBMs), which use textual rationales as intermediates for vision classification, justifying a single known label per input. While our Language Bottleneck Models (LBMs) adopt this language bottleneck concept, they differ fundamentally. Unlike XBMs’ instance-specific, task-specific rationales, LBM summaries aim to capture an implicit knowledge states—as emphasized by our inverse problem formulation—and generalize to future, unknown questions. These requirements necessitate holistic, adaptable summaries rather than XBMs’ task-specific rationales, leading us to introduce Language Bottleneck Models (LBMs) as a distinct framework with broader applicability.
Contrast with Text Summarization Models
Recent text summarization models such as BART lewis2019bart , T5 raffel2020exploring , and PEGASUS zhang2020pegasus effectively produce concise summaries based on explicitly available textual content. However, these approaches differ fundamentally from knowledge tracing, where summaries must infer implicit student knowledge states not directly observable in the input. Standard summarization metrics (e.g., ROUGE, BLEU) rely on explicit reference summaries, making them unsuitable for evaluating latent knowledge inference tasks. In contrast, our Language Bottleneck Models generate textual summaries of the student’s implicit knowledge state, optimized for predictive accuracy on downstream questions rather than syntactic overlap with the observed input.
5 Experiments
We evaluate the effectiveness of our Language Bottleneck Models (LBM) in the context of Knowledge Tracing (KT), both with a controlled synthetic KT task involving basic arithmetic operations, and on real-world knowledge tracing datasets.
5.1 Experimental Setup
Datasets We evaluate on one synthetic and one real-world source. Synthetic arithmetic simulates learners answering addition, subtraction, multiplication, and division items; each student is assigned mastered skills, unmastered skills, and systematic misconceptions. The Eedi dataset wang2020diagnostic contains millions of question–answer pairs from UK secondary-school mathematics. More details about these datasets are provided in the Appendix.
Models The main experiments use GPT-4o and GPT-4o-mini, together with open-source Qwen 2.5 (3B, 7B) and Gemma 3 (12B, 30B) checkpoints. Unless noted, we run the instruct variants of each open-source model; hyper-parameters and prompt templates are provided in the Appendix.
Baselines We compare LBMs against (i) four standard knowledge-tracing methods—DKT piech_deep_2015 , DKVMN zhang2017dynamic , SAKT pandey2019self , and AKT ghosh2020context —implemented with the pyKT library liupykt2022 ; and (ii) direct prompting, where each LLM predicts answers from the full student history without a bottleneck. Training details for all baselines are given in Appendix.
5.2 Results
5.2.1 LBM vs Direct Prompting
Table 1: Accuracy results for Direct and LBM methods. Bold indicates LBM accuracy is no more than 2% below the Direct baseline.
synthetic eedi Direct LBM $\Delta$ Direct LBM $\Delta$ Qwen2.5-3B 0.525 0.627 0.102 0.347 0.372 0.025 Qwen2.5-7B 0.655 0.646 -0.009 0.647 0.583 -0.065 gemma-3-12b 0.790 0.791 0.001 0.580 0.625 0.045 gemma-3-27b 0.785 0.794 0.009 0.670 0.650 -0.020 gpt-4o-mini 0.795 0.782 -0.013 0.655 0.645 -0.010 gpt-4o 0.840 0.825 -0.015 0.703 0.682 -0.020
Figure 4: Evolution of accuracy with bottleneck length.
Table 1 compares the accuracy of LBM models to direct prompting variants across different LLMs. Despite the information bottleneck enforced by LBMs, they remain within $2\%$ of the accuracy obtained via direct prompting. Figure 4 demonstrates that this gap decreases for longer bottlenecks, highlighting a trade-off between conciseness and predictive accuracy.
Encoder/decoder variants
We evaluate LBMs with different models as encoders and decoders (results shown in the Appendix). Using the strongest model (gpt-4o) as encoder with other models from Table 1 as decoders yields accuracies $5-10\%$ higher than when the stronger model is used as the decoder. This confirms our hypothesis that extracting relevant information for the summary (encoding) is more challenging than predicting future answers given a summary (decoding).
5.2.2 LBM vs Knowledge Tracing Methods
Figure 5 compares the performance of LBM models to traditional Knowledge Tracing methods on the synthetic dataset. LBMs with strong LLMs achieve comparable accuracy to KT methods with significantly less training data. Since KT methods rely on statistical patterns, they require substantially more observations before reaching strong predictive power, while LBMs demonstrate effective generalization with minimal examples.
Figure 5: Accuracy of LBM vs Knowledge Tracing models on the Synthetic dataset. The x axis represents the total number of question seen by the model, ie the product of number of questions observed per student $T$ and the number of students available for training $N$ (note that N=1 for LBM methods). The y axis is the average accuracy 200 test student and 4 new questions to predict per student. The pre-trained models indicated for LBM are used as both Encoder and Decoder, with chain-of-thought prompting. Error bars report the standard error calculated across 10 runs for N=10 students and 3 runs otherwise.
5.2.3 Training LBM Encoders
As discussed in Section X, downstream accuracy serves as a reward signal to train encoders to produce increasingly accurate summaries. We demonstrate this by training a Gemma3-12B encoder with GRPO alongside a frozen Gemma3-27B decoder. Figure 6 shows the encoder progressively improving summary quality, quickly outperforming gpt-4o. The figure illustrates this progression through an example student who mastered all constructs except subtraction and fails any multiplication involving 9. The initial summary contains inaccuracies and misses this systematic misconception, while the final summary successfully captures the student’s complete knowledge state.
Figure 6: Training the encoder with the decoding accuracy on the synthetic dataset. Evolution of the test accuracy as a Gemma3-12B encoder is trained as described in Section 3.3, with a fixed Gemma3-27B decoder. Trained on $800$ training students and tested on $200$ students, with $|\mathcal{X}|=50$ questions per input trajectory and $|\mathcal{Y}|=20$ questions to predict per student. The bottom row show the evolution of the generated summary over the course of training for an example student. Text is colored green (exact), orange (approximate), or red (false) based on ground-truth.
Accuracy gains from training are stronger on harder cases
When stratifying students by difficulty (measured by number of misconceptions), we find training yields larger accuracy improvements for complex cases (see table in the Appendix). The model shows 10-11% increased accuracy relative to gpt-4o for students with 2 or 3+ misconceptions, compared to just 5-7% for students with 0-1 misconceptions. This suggests the encoder becomes particularly better at handling complex knowledge states.
5.2.4 Steering LBM behavior
We demonstrate multiple steering strategies described in Section 3.4 (details in Appendix). We find that providing explicit misconception information during encoder training produces significantly stronger learning effects than adding the same information at the decoder stage. This suggests the encoder leverages this additional context to better interpret patterns in student responses. We also show the encoder can be steered during training to explicitly mention misconceptions through reward signals, and illustrate how supplementing the bottleneck with information not present in the input can significantly improve decoder accuracy.

### v2

@@ 3.3.1 Encoder
Optimization via GRPO We optimize the encoder $f_{\theta}$ with Group Relative Policy Optimization (GRPO) (Shao et al., 2024). For each input $\mathcal{X}_{\text{enc}}$, the encoder generates $G$ candidate summaries $\{\tilde{\mathcal{S}}^{1},\ldots,\tilde{\mathcal{S}}^{G}\}$, each receiving reward $r_{i}=R(\tilde{\mathcal{S}}^{i};g)$. We then compute group-relative advantage $\hat{A}_{i}=\frac{r_{i}-\text{avg}(\textbf{r})}{\text{std}(\textbf{r})}$ and train with the DAPO loss (Yu et al., 2025):
$\mathcal{L}(\theta)=-\frac{1}{\sum_{i=1}^{G}|\tilde{\mathcal{S}}^{i}|}\sum_{i=1}^{G}\sum_{t=1}^{|\tilde{\mathcal{S}}^{i}|}\biggr[\frac{\pi_{\theta}(\tilde{\mathcal{S}}^{i}_{t}|\mathcal{X}_{\text{enc}},\tilde{\mathcal{S}}^{i}_{<t})}{[\pi_{\theta}(\tilde{\mathcal{S}}^{i}_{t}|\mathcal{X}_{\text{enc}},\tilde{\mathcal{S}}^{i}_{<t})]_{\text{no grad}}}\hat{A}_{i}\\ -\beta\mathbb{D}_{\mathrm{KL}}[\pi_{\theta}||\pi_{\text{ref}}]\biggr]$ (2)
where $\beta$ is a hyperparameter and $\pi_{\text{ref}}$ is the original model.
3.3.2 Decoder
The decoder $g_{\phi}$ is trained via standard supervised fine-tuning (SFT). We generate knowledge state summaries $\tilde{\mathcal{S}}$ for a set of training students and pair them with ground-truth correctness labels for target questions $q\in\mathcal{Y}$ . The model is optimized to predict the correct binary outcome (e.g., "Yes" or "No") conditioned on the summary $\tilde{\mathcal{S}}$ and query question.
3.4 Steerability of the Estimated Knowledge State
@@
The natural-language summaries generated by LBMs allow for various human-model interactions (detailed in Appendix D). (1) Prompt engineering the encoder. Since the encoder $f_{\theta}$ is itself an LLM, its behavior can be shaped through prompt design, such as system instructions or in-context examples (Brown et al., 2020). (2) Steering via reward signals. Rewards to steer the encoder towards human preferences can be incorporated through the $\Omega(S)$ term in Eq. 1. (3) Augmenting with student-specific information. Educators can supplement the model with additional knowledge not present in observed data—either by augmenting encoder inputs or by editing the generated summary before decoding. This enables integration of recent classroom observations or specific misconceptions identified through in-person interactions.
4 Experiments
Datasets
We evaluate LBMs and baseline models on a synthetic arithmetic benchmark and two real-world datasets (Table 4). Our Synthetic dataset (Appendix  B.4.1) simulates learners answering addition, subtraction, multiplication, and division questions; each student is assigned mastered skills, unmastered skills, and systematic misconceptions. We filter Eedi and XES3G5M for single-session trajectories (max inter-question gaps of 3 and 10 minutes respectively) with at least 40 and 34 interactions respectively to approximate the static knowledge state assumption common in CD (Wang et al., 2024). We evaluate XES3G5M in both Chinese and English using translations from Ozyurt et al. (2024) (see Appendix B.4).
Table 1: Overview of datasets. AVG#log and STD#log>1 are defined following  Wang et al. (2022) as respectively the average number of logs per student per KC, and the mean standard deviation of score per student and per KC.
Dataset Synthetic Eedi (Filt.) XES3G5M (Filt.)
#Students 2,000 623 996
#Questions 5,000 3,401 3,221
#KCs 4 1,284 803
#Logs 420,000 28,947 57,788
#Logs/stud. 210 $\geq$40 $\geq$34
[…]
Figure 3: Case study: comparing CD and LBM knowledge states. Given a student from the Synthetic dataset, we compare proficiency estimates across knowledge concepts (KCs) obtained from a trained NeuralCDM model to the text-based knowledge state generated by a trained LBM model.
Models We evaluate LBMs across LLM backbones of different sizes and capabilities, both open-source (Qwen 2.5 3B and 7B (Team, 2024), Gemma 3 12B and 27B (Team, 2025)), and closed-source (GPT-4o-mini, GPT-4o and GPT-5 (Achiam et al., 2023)). We train one decoder model by finetuning a Gemma3-12B model with LoRA using on a GPT-5 encoder (labeled gpt-5+gem-12b* in the results, see details in § B.2). Unless noted otherwise, the same backbone LLM is used for both the encoder and decoder, and prompt all models to provide their response directly without chain-of-thought. We run GPT-5 with reasoning_effort=minimal configuration. Hyper-parameters and prompt templates are provided in Appendix B.
Baselines We compare LBMs against 9 Knowledge Tracing methods: DKT (Piech et al., 2015), DKVMN (Zhang et al., 2017), SAKT (Pandey and Karypis, 2019), AKT (Ghosh et al., 2020), Deep IRT (Yeung, 2019), SAINT (Choi et al., 2020), SimpleKT (Liu et al., 2023a), QIKT (Chen et al., 2023), GKT (Nakagawa et al., 2019)) implemented with the pyKT library (Liu et al., 2022) and 5 Cognitive Diagnosis methods (IRT (Rasch, 1993), MIRT (Reckase, 2006), DINA (Junker and Sijtsma, 2001), KaNCD and NeuralCDM (Wang et al., 2022)) implemented with the EduCDM library (bigdata-ustc, 2021). We also run each LLM via direct prompting, where the LLM predicts answers from the full interaction history without a bottleneck. Training details for all baselines are given in Appendix B.
Systematic evaluation of knowledge state summaries We systematically evaluate the quality of the knowledge state summaries produced by different encoder models on the Synthetic dataset using LLM-as-a-judge (Li et al., 2025). Specifically we prompt a GPT-5 model to compare each summary to the corresponding ground-truth knowledge state across the following dimensions (see Appendix C for full details and prompts):
  • Global score: overall alignment with the true state,
  • Construct accuracy: correctly assessed constructs,
  • Misconception detection: correctly identified student misconceptions,
  • Misconception false positives: misconceptions mentioned but absent in the ground truth,
  • Specificity: precision vs. vagueness of the description,
  • Confidence calibration: degree of over- or under-confidence.
Figure 4: Systematic evaluation of the summaries produced by different encoder LLMs over 200 test students from the Synthetic dataset: (Left) Average construct mastery accuracy across constructs (%) and average overall score (1-5 scale); (Middle) Average misconception detection rate (%) and average number of misconception false positives per summary (count); (Right) Confidence Index (1-3, 1=under-confident, 2=appropriately calibrated, 3=over-confident) and Specificity Index (1-3, higher is more specific). Error bars represent the standard error (N=200).
4.1 Qualitative Insights
4.1.1 Case-study: comparing CD and LBM knowledge state representations
A key advantage of LBMs over CD models is the ability to capture nuanced insights about the student knowledge state, such as misconceptions. We illustrate this with a case-study Figure 3. We train a state-of-the-art CD model (NeuralCDM) trained on our Synthetic dataset, which achieves strong predictive performance (AUC: 0.96, Acc: 0.90 on the test set). From this model, we extract proficiency estimates across knowledge concepts using the learned student embedding vector, producing a typical CD diagnostic report. We then compare this to the knowledge state summary generated for the same student by an LBM (trained Gemma-12B encoder with frozen Gemma-27B decoder, as in §4.2).
While NeuralCDM reliably captures general KC proficiency, its estimates are influenced by misconceptions without explicitly identifying them. In contrast, LBMs capture overall proficiency and uncover specific misconceptions (e.g., errors with negative numbers or with operand-6). This ability to provide nuanced, qualitative insights into student knowledge states sets LBMs apart from CD methods.
4.1.2 Systematic evaluation of knowledge summaries
To further assess the interpretability of the knowledge states produced by LBMs, we systematically compare their summaries to ground truth knowledge states from the Synthetic dataset using GPT-5 as LLM-as-a-judge (Li et al., 2025). The results are shown in Figure 4. Figure 4a shows the overall alignment with the ground truth summary (score from 1=’mostly incorrect’ to 5=’strongly aligned’), and the construct mastery accuracy. Model capability correlates with performance on both metrics, with GPT-5 achieving the highest scores. Figure 4b plots the misconception-detection rate against the misconception false positive rate. Qwen2.5 and GPT-4o/4o-mini exhibit lower hallucination but also lower detection; Gemma-3 detects more misconceptions but at the cost of more false positives; GPT-5 outperforms all models on both dimensions. Figure 4c evaluates summary specificity and confidence calibration. Gemma-3 models produce more specific summaries than Qwen2.5, but both families tend to be over-confident relative to GPT models, with GPT-5 displaying the best overall specificity and calibration.
Refer to captionA
Refer to captionB
Figure 5: Training the encoder on Synthetic dataset. (A) Evolution of the test accuracy during training of a Gemma3-12B encoder, with a frozen Gemma3-27B decoder. Trained on $800$ students and tested on $200$ students, with $|\mathcal{X}|=50$ questions per input trajectory and $|\mathcal{Y}|=20$ questions to predict per student. Bottom row shows the evolution of generated summaries over the course of training for an example student. Text is colored green (exact), orange (approximate), or red (false) based on ground-truth. (B) Systematic evaluation of generated summaries over encoder training. Evaluation axes are the same as Figure 4(a) and (c). Confidence intervals show the standard error (N=200).
4.2 Training the Encoder
As outlined in §3.3.1, downstream accuracy can serve as a reward signal to train an encoder to produce increasingly accurate summaries. We demonstrate this by training a Gemma3-12B encoder using a frozen Gemma3-27B decoder on 800 students. We set the reward as the decoder accuracy across $|\mathcal{Y}|=20$ unseen questions $R(\tilde{\mathcal{S}};g)=\texttt{Acc}\big(\tilde{\mathcal{Y}},\mathcal{Y}\big)$, train with a LoRA adapter (Hu et al., 2022) and evaluate on 200 unseen test students.
Effect on accuracy Figure 5 shows the encoder progressively improving summary quality and quickly outperforming GPT-4o. The figure illustrates this through an example student who mastered all constructs except subtraction and fails any multiplication involving 9. The initial summary contains inaccuracies and misses this systematic misconception, while the final summary successfully captures the student’s complete knowledge state. Stratifying by knowledge state complexity, we observe larger gains for more complex cases (Appendix A.3.1).
Qualitative evolution of summaries We use the same LLM-as-a-judge approach as in the previous section to systematically evaluate the generated summaries over training. Figure 5 shows that the summaries gradually better capture concept mastery and align more with the ground truth knowledge state over training. Moreover, as summaries also become more specific and better calibrated. Misconception detection and false positive rates do not significantly change over the course of training (Appendix A.3.2).
4.3 LBM vs Knowledge Tracing Methods
We present comprehensive results comparing LBMs to baseline methods across all datasets (detailed results in Table A.2 in the Appendix). Here we highlight key findings and insights from these experiments.
Refer to captionA
Refer to captionB
Figure 6: (Left) Accuracy of LBM vs the top-2 KT and CD models across datasets. Models are grouped by LBMs (blue), KT models (orange) and CD models (red). Top 2 KT and CD models are selected based on average accuracy across all three datasets. LBMs are evaluated zero-shot except gpt-5+gem-12b* where the Gemma3-12b decoder is trained. Full results for other models are available Table A.2 in the Appendix. (Right) Accuracy of the best LBM, KT and CD models on the Synthetic dataset. The x axis shows the total number of question seen by the model, i.e. "Traj. Length" x "# Students". Note that $\text{Students}=1$ for LBMs as they are evaluated zero-shot. Results are averaged across N=10 runs for KT and CD models and N=3 for LBMs, with error bars showing the standard error.
Performance Figure 6 compares zero-shot and trained LBMs against the top KT and CD models (full results in Table A.2). Zero-shot LBMs with powerful backbones (GPT-4o, GPT-5) rival the performance of the best baselines across all datasets. Training the decoder further improves accuracy: a fine-tuned Gemma-12B decoder outperforms the zero-shot GPT-5 decoder when both use the same GPT-5 encoder, and further closes the gap in accuracy with KT/CD models. Notably, this "GPT-5 Encoder+trained Gemma3-12b Decoder" configuration ranks first in AUC among all KT/CD models on Eedi and third on XES3G5M-English. (Table A.2).
Sample efficiency Figure 6B compares the performance of the best zero-shot LBM model (GPT-5) to the best CD and KT models (KaNCD and AKT) on the Synthetic dataset. LBMs with a GPT-5 backbone achieves comparable accuracy to KT/CD methods with significantly less training data. Since CD/KT methods rely on statistical patterns, they require substantially more observations before reaching strong predictive power, while LBMs demonstrate strong zero-shot performance.
4.4 Steering LBM behavior
We demonstrate multiple steering strategies described in §3.4. Providing explicit misconception information during encoder training produces substantially stronger learning effects than adding the same information at the decoder stage (Appendix A.4.1), suggesting that the encoder uses this additional context to better interpret patterns in student responses. We also show the encoder can be steered during training to explicitly mention misconceptions through reward signals (Appendix A.4.2), and illustrate how supplementing the summary with information not present in the input can significantly improve decoder accuracy (Appendix A.4.3).
4.5 Ablation Experiments
How much information is lost by the bottleneck? Table 2 compares the accuracy of LBM models to directly predicting new questions from the observed student data. Despite the information bottleneck, LBM accuracy typically remains within $2\%$ of direct prediction—and often surpasses it. Figure 7 shows that this gap decreases for longer bottleneck token limits, highlighting a trade-off between conciseness and predictive accuracy.
Which of the encoder or decoder is most critical for LBMs? We evaluate LBMs with different encoder–decoder pairings (Appendix A.5.1). Using a strong model (GPT-4o) as the encoder with weaker models as decoders yields accuracies $5-10\%$ higher than when the stronger model is used as the decoder. This confirms our hypothesis that extracting relevant information for the summary (encoding) is more challenging than predicting future answers given a summary (decoding).
Do LBMs require knowledge concept information? Table LABEL:app:tab:ablation compares LBMs with and without KC information in the input prompt on the Synthetic and EEDI (Filtered) datasets. Performance does not significantly change, demonstrating that LBMs do not fundamentally require KC information.
Figure 7: Evolution of LBM accuracy with bottleneck length (GPT-4o backbone).
Table 2: Accuracy results for Direct and LBM methods. Bold indicates models for which the LBM accuracy is no more than $2\%$ below the Direct baseline (Welch’s t-test, details in Appendix B.5.1).
synthetic eedi Direct LBM $\Delta$ Direct LBM $\Delta$ Qwen2.5-3B .56 ± .00 .61 ± .01 +.06 .35 ± .00 .38 ± .00 +.03 Qwen2.5-7B .64 ± .00 .65 ± .01 +.01 .65 ± .00 .58 ± .01 -.07 gemma-3-12b .63 ± .00 .79 ± .00 +.17 .58 ± .00 .62 ± .01 +.04 gemma-3-27b .62 ± .00 .78 ± .01 +.16 .67 ± .00 .65 ± .01 -.02 gpt-4o-mini .81 ± .01 .78 ± .01 -.03 .67 ± .01 .61 ± .02 -.05 gpt-4o .85 ± .01 .80 ± .01 -.05 .66 ± .00 .66 ± .02 +.00 gpt-5 .87 ± .00 .87 ± .01 +.00 .71 ± .02 .68 ± .01 -.03
@@ 5 Related Work
We review related works from the Cognitive Diagnosis and Knowledge Tracing literature, as well as concept bottleneck models. See Appendix F for an extended review of related works, and Table F1 for a high-level comparison of LBMs with CD and KT.
Cognitive Diagnosis Cognitive Diagnosis Models (CDMs) infer student knowledge states from observed responses. Classical approaches include Item Response Theory (IRT) and Multidimensional IRT which measure continuous proficiency scores (Rasch, 1993; Reckase, 2006), and the DINA model and its variants which estimate binary mastery of knowledge concepts (De La Torre, 2009). Recent deep learning variants like NeuralCDM (Wang et al., 2022) and RCD (Gao et al., 2021) use neural networks and graph architectures to model complex relationships between students, questions, and knowledge concepts. However, these models provide only quantitative skill mastery estimates over knowledge concepts.
Knowledge Tracing Knowledge Tracing methods model student learning to predict future performance. Deep learning approaches like DKT (Piech et al., 2015), DKVMN (Zhang et al., 2017) and AKT (Ghosh et al., 2020) employ neural architectures. Despite strong predictive performance, these models represent knowledge as abstract latent vectors, limiting interpretability. Several recent works have proposed more interpretable KT. Early Bayesian approaches (Corbett and Anderson, 1994; Käser et al., 2017) and recent probabilistic models (Minn et al., 2022; Zhou et al., 2024) learn interpretable latent states representing student proficiency over knowledge concepts, while IRT-based methods (Yeung, 2019; Chen et al., 2023) combine deep learning with item response theory for meaningful latent representations. Other works improve interpretability through learned question relationships (Tong et al., 2022), explainable subsequences (Li et al., 2023), or option tracing (Ghosh et al., 2021). However, these approaches remain fundamentally constrained to quantitative concept proficiency estimation or require post-hoc interpretability. Finally, recent LLM-based approaches have been proposed for knowledge tracing tasks (Li et al., 2024a; Wang et al., 2025), but they generally remain opaque, directly predicting future answers without enforcing interpretable intermediate representations.
@@
Concept Bottleneck Models Concept Bottleneck Models (CBMs) (Koh et al., 2020) improve interpretability by using human-understandable concept activations as intermediates between inputs and predictions. However, CBMs typically rely on finite predefined concept sets, limiting applicability to complex tasks like knowledge tracing. Recently, Explanation Bottleneck Models (XBMs) (Yamaguchi and Nishida, 2024) use textual rationales as intermediates for vision classification. While our Language Bottleneck Models (LBMs) adopt this language bottleneck concept, they differ fundamentally: unlike XBMs’ instance-specific rationales, LBM summaries capture implicit knowledge states that generalize to future, unknown questions, requiring holistic, adaptable summaries rather than task-specific rationales.